Translate the following statement into an algebraic equation. Let x represent the number. 7 less than 10 times a number is 9 times that same number.



Answer by tutorcecilia(2152)

10x-7=9x
10x-9x-7=9x-9x [solve for x]
x-7=0
x=7
.
check by plugging x=7 back into the original equation:
10x-7-9x
10(7)-7=9(7)
70-7=63
63=63 [checks out]